OpenStudy (mikezack123):

(-1, 3) and perpendicular to y = -1/3x + 7.

OpenStudy (mikezack123):

y = 3x - 12 y = 3x + 6 y = 3x - 6 y = 3x

OpenStudy (anonymous):

You need an equation that passes through (-1,3)?

OpenStudy (mikezack123):

@AlwaystheBookworm yes

OpenStudy (mikezack123):

and that is perpendicular to that equation above.

OpenStudy (mathstudent55):

What is the slope of the given line?

OpenStudy (anonymous):

Oh, well you could simply just plug it into the point-slope formula.

OpenStudy (mikezack123):

@AlwaystheBookworm ... OH yeah! lol

OpenStudy (anonymous):

@MikeZack123 lol if you're still having trouble with it, let me know and I'd be happy to help.

OpenStudy (mikezack123):

@AlwaystheBookworm i got B... is that right?

OpenStudy (anonymous):

remember that to get the slope of the perpendicular line you need to inverse opposite. so inverse the fraction or number into a fraction and put the opposite sign

OpenStudy (mikezack123):

@melody16 i dont get what youre saying... the answer is not b?

OpenStudy (anonymous):

She means that once you solve the equation regularly, you have to find the inverse.

OpenStudy (anonymous):

I am just saying, for future reference. alwaysthebookworm explained what i meant

OpenStudy (mikezack123):

... im confused i dont get why you are saying that, you didnt have to solve the equation in this problem so.....

OpenStudy (whpalmer4):

(-1, 3) and perpendicular to y = -1/3x + 7. Okay, the slope of the original line is -1/3 because it is in slope-intercept form: \[y = mx+b\] where \(m\) is the slope. We need to find the slope of a perpendicular line. If you have a slope \(m\) and need to find a perpendicular slope \(p\), you can find it by evaluating \(p = -1/m\). Our perpendicular slope will be \(p = -1/(-1/3) = -1 * (3/-1) = -3/-1 = 3\) As a check, the product of two perpendicular slopes is -1 3 * (-1/3) = -3/3 = -1. So far so good. we know our perpendicular line has to go through point (-1,3) and have slope 3. We'll use the point-slope equation for a line with known slope going through a point \((x_0,y_0)\): \[y-y_0 = m(x-x_0)\] (note the \(m\) here is going to be our perpendicular slope, not the slope from the original line) \[y -3 = 3(x-(-1))\]rearrange to get in slope-intercept form \[y = 3x +6\]

OpenStudy (mikezack123):

thanks to all of you! :)

OpenStudy (mikezack123):

i wish i could give you ALL medals :(